Evaluate the limit limh→0
2(−3+h)2−18
h

limh→02(−3+h)2−18h
1
Expert's answer
2019-04-01T11:42:47-0400

"\lim\limits_{h\to 0}(2(-3+h)^2-18h)= \lim\limits_{h\to 0}(2(9-6h+h^2)-18h)="

"=\lim\limits_{h\to 0}(18-12h+2h^2-18h)
=\lim\limits_{h\to 0}(18-30h+2h^2)=" "\{30h \to 0 \quad as\quad h\to 0 \quad and \quad 2h^2\to 0 \quad as\quad h\to 0,\quad \\hence \quad(18-30h+2h^2)\to 18\quad as\quad h\to 0\}="

=18.